1. You need to prepare 10.0 liters of an acid aqueous solution with a pH of 1.25. Calculate the mass of sulfuric acid that you need (in grams).
Answer #1

Ans :

[H+ ] = 10-pH

putting value of pH

[H+ ] = 10-1.25

[H+ ] = 0.0562 M

M = mol solute / volume (L)

0.0562 M = n / 10.0 L

n = 0.562 mol

Each mol H2SO4 gives 2 mol H+ ions , so the number of mol of H2SO4 here will be : 0.562 mol / 2 = 0.281 mol

Mass of sulfuric acid = mol x molar mass

= 0.281 mol x 98.079 g/mol

= 27.6 grams